FOUR TUTORIALS ON IMAGES: NOVICE → EXPERT April 19, 2015

Carl Heiles

Images are important. We have IDL procedures that help you do images and we have prepared four tutorials on images that explain things, including our procedures. These tutorials take you from the know-nothing novice to the overconfident expert. Like a four-course French dinner, you ingest them in order of presentation. They’re bite-sized; don’t gobble, finish each course before trying the next. Here they are:

1. You need to know plotting to do images, e.g. if you want to label them with axes and annotations. For this you need bpidl.ps, bpidl.pdf: BPIDL—BASIC PLOTTING IN IDL... It covers…

(a) Basic plotting, overplotting

(b) Sampling the plot with the cursor (we like Tim Robishaw’s trc.pro

(c) Annotating plots; Greek letters, etc., using textoidl

(d) Arrays of plots

(e) Plottine up to five variables simultaneously using plotting symbols, color, size

(f) Selecting groups of points in a plot with the cursor

(g) Making good-looking PostScript files

2. You need to understand basic concepts and to learn how to display and manipulate images on your computer screen. Enter our introductory tutorial ididl.ps, ididl.pdf: Image Display and Manipulation in IDL (IDIDL) ... Images with a Single 256-Entry Color Table. It covers…

(a) The X-window and your computer screen; pixels, bytes.

(b) TrueColor, PseudoColor, the color table, combined color, decomposed color

(c) The Visual Class: Dynamic vs. static colortables

(d) Displaying the image with IDL’s tv and with Tim Robishaw’s display.pro .

(e) Image manipulation, processing, contrast enhancement.

(f) Colorbars: one-d (intensity only) and two-d (intensity and color)

3. Ordinary images display a single quantity, often surface brightness; we call these one-dimensional images. However, in a single image you can use color to display two or three quantities simultaneously. We call these two-dimensional and three-dimensional images, respectively., so appro-

priately enough this final tutorial (newdimensionalcolor.ps, newdimensionalcolor.pdf

is entitled 1d2d3d: One, Two, and Three Dimensional Color Images… (a) One-dimensional images; contrast; colortables; the colorbar

(b) Non-gray one-dimensional color images: choosing the colortable.

(c) Two-dimensional color images: intensity combined with a non-gray colortable; using nonsaturated colors to make perceived intensity orthogonal to color.

(d) Three-dimensional color images: the orthogonality of red, green, and blue.

(e) Colorbars for one-, two-, and three-dimensional color images.

(f) Appendices, which tell how to make these images in IDL.

====================================================

4. The first course dealt with the bare image. That’s like having an hors d’oeuvre without wine: usually you want to annotate the image by labelling with axes or putting arrows or labels on the image itself. Moreover, you should indicate what a particular color or intensity on the image means in terms of physical units (such as surface brightness), for which you use a colorbar. So the next course is IACIDL: Image Annotation and Colorbars in IDL It covers…

(a) How to insert your image in the window and leave appropriate space around the border for axes and/or a colorbar.

(b) How to make labelled axes and insert comments on images.

(c) How to insert a colorbar.

(d) How to scale an arbitrary-sized image so that it fits on your computer screen.

(e) Important basic details for making X-windows and PostScript images.

5. The first two courses concentrate on images for the computer screen. That’s like cooking for yourself: it’s fine for you, but you usually want to present your final, beautifully handcrafted image to the world, either on paper or in (God Forbid!) PowerPoint. The first step is making a PostScript image, so the next course is entitled PSIDL: PostScript Files in IDL – Making, Printing, and PowerPointing. Note this works for graphs as well as images! Your PostScript images can be either objects of beauty or… not. We much prefer the former and this tutorial tells all about it…

(a) Greek letters and Fonts in X and PS: psopen and textoidl.

(b) Plotting directly to PostScript.

(c) Printing PostScript images; PowerPointing PostScript images.

(d) Lots of appendices with details you might wwant to know, including how to make quick and dirty PostScript directly from your computer screen image.